OpenStudy (anonymous):

Idenitify the 19th term of the arithmetic sequence 5,16,27... ?

OpenStudy (anonymous):

a=5 d=11 so 19th term = a+(n-1)d= 5+18*11=5+198=203

OpenStudy (anonymous):

thanks